OpenStudy (anonymous):

What does (3 + deltax)squared equal? Please show workings

OpenStudy (anonymous):

3+ deltax simply means theres a slight increase in the value of 3 but its unknown.

OpenStudy (anonymous):

when its squared u will get 9 + delta square x + 6deltax

OpenStudy (anonymous):

But why isn't it just 9 + deltaxsquared?

OpenStudy (anonymous):

because of property in algebra (a+b)^2= a^2+b^2+ 2ab

OpenStudy (anonymous):

have you learNT this formula ?

OpenStudy (anonymous):

no. doesn it have a name that I can google for tutorials?

OpenStudy (anonymous):

dear, in what class are you in?

OpenStudy (anonymous):

I am an adult who has forgotten so much and is trying to learn by self. I don't know the name of stuff.

OpenStudy (anonymous):

Laws of Exponents (am)(an) = am+n (ab)m = ambm (am)n = amn Fractional Exponents a0 = 1 (am)/(an) = am-n a-m= 1/(am) Quadratic Formula In an equation like ax2 + bx + c = 0 You can solve for x using the Quadratic Formula: Quadratic Formulas Binomial Theorem (a + b)1 = a + b (a + b)2 = a2 + 2ab + b2 (a + b)3 = a3 + 3a2b + 3ab2 + b3 (a + b)4 = a4 + 4a3b + 6a2b2 + 4ab3 + b4 ...and so on... Difference of Squares a2 - b2 = (a - b)(a + b) Rules of Zero 0/x = 0 where x is not equal to 0. a0 = 1 0a = 0 a*0 = 0 a/0 is undefined (you can't do it)

OpenStudy (anonymous):

WOW thanks!

OpenStudy (anonymous):

cut that// i will give you the link.. http://www.freemathhelp.com/algebra-formulas.html

OpenStudy (anonymous):

Thankyou so much!!!!